OpenStudy (ytry):

The graph of a quadratic function is a parabola that opens down and has a vertex of (4, 3). Which of the following could be the function? A. y = -x^2 +4x +3 B. y = -x^2 +8x -13 C. y = x^2 -4x + 3 D. y = x^2 -8x -13

OpenStudy (cwrw238):

hint: if it opens down then it will have - before the x^2 writing it in vertex form will then show you which one it is

OpenStudy (cwrw238):

vertex form is y = (x+a)^2 + b

OpenStudy (ytry):

I think it is B but I am not sure

OpenStudy (ytry):

NM it is A

OpenStudy (cwrw238):

you sure its A?

OpenStudy (ytry):

From what you said I think it is atleast

OpenStudy (ytry):

IF vertex form is what you said then it should fit

OpenStudy (ytry):

Please do correct me if I am wrong though, I am honestly struggleing in algebra

OpenStudy (cwrw238):

y = -x^2 +8x -13 = - (x^2 - 8x + 13) = -[ (x - 4)^2 -3)] at vertex x = 4 because that gives maximum value of the function now you can work out what y is at vertex

OpenStudy (cwrw238):

ok? y = 3 so vertex is (4,3) so its B

OpenStudy (ytry):

Thanks for your help

OpenStudy (cwrw238):

yw